OpenStudy (anonymous):

Lesson: Estimating with Fractions and Mixed Numbers Content Includes: -Estimating each sum or difference. -Estimating each product or quotient

OpenStudy (anonymous):

Estimate Each Sum or Difference: 1. \[\frac{ 1 }{ 6 } +\frac{ 5 }{ 8 } \approx \] 2. \[\frac{ 7 }{ 8 } +\frac{ 1 }{ 16 } \approx \] 3. \[\frac{ 1 }{ 10 } +\frac{ 5 }{ 6 } \approx \]